OpenStudy (anonymous):

Last question! Use the Pythagorean Theorem to determine if a triangle with side lengths of 8 inches, 20 inches, and 15 inches is a right triangle. Answer= Yes? or No?

OpenStudy (helder_edwin):

the pythagorean theorem says \[ \large a^2+b^2=c^2 \] c is the hypothenuse and is the longest side of the triangle then you tell me a=? b=? c=?

OpenStudy (anonymous):

a=8 b=15 c=20 ???

OpenStudy (helder_edwin):

yes now square them and plug them into the equation

OpenStudy (anonymous):

y=8^2+15^2=20^2 y=64+225=400 ?

OpenStudy (helder_edwin):

a^2=64 b^2=225 c^2=400 tell me is it true that \[ \large a^2+b^2=c^2 \]

OpenStudy (anonymous):

no.

OpenStudy (helder_edwin):

so is your triangle recrtangle?

OpenStudy (anonymous):

no?

OpenStudy (helder_edwin):

yes! great job!

OpenStudy (anonymous):

thank you!

OpenStudy (helder_edwin):

u r welcome